# Magneto-optical trap loading with an effusive oven in    a large optical access experiment

### v1

@@ B. The experiment and comparison to
In Fig. 2, we compare the 3D MOT loading diagrams obtained analytically, experimentally, and numerically. The parameter space spans over different field gradients and top beam detunings. The shapes enveloping significant atom numbers are observed to take up roughly similar areas in all three cases, with the highest atom numbers (absolute values discussed below) obtained at the same data point locations in the experiment and simulation (dotted circle center, $48$ G/cm gradient at $-10$ MHz top beam detuning), while the analytical model prediction for the highest number is within the point’s neighborhood (cross, $44.8$ G/cm gradient at $-7.2$ MHz top beam detuning). The analytical model (Sec. III.A) can be used to explain the different features in these diagrams as follows.
The dark area near the line where the axial beams are equally detuned ($-46$ MHz; vertical dashed line) is com-
@@
posed of three parts. Just to the left of this line, the transfer occurs in the opposite direction (larger top beam detuning). At this line, the transfer is obviously halted. Just to the right, the transfer is slow enough so the atoms disperse in significant numbers before reaching the 3D MOT region, with the  higher  gradients  resulting in the dark area extending further towards resonance as the radial capture distance $d_{\textsf{c}}$ becomes smaller.
@@
The dark area on the right side of the diagrams (delineated by the slanted dashed line) is, on the other hand, obtained due to the axial equilibrium position $d_{\textsf{e}}$ being far enough below the radial beams center so the confinement is significantly diminished, or even non-existent due to the radial beams not encompassing $d_{\textsf{e}}$. The boundary appears linear as $d_{\textsf{e}}$ is proportional to the axial beam detuning difference while inversely proportional to the gradient.
@@
Regarding the bright isle with the highest atom numbers, its location depends on the MOT confinement and the dark-state loss. The former expresses itself in the capture probability (see the error functions in Eq. 1a) and the relative positioning of the radial beams (see the Gaussian in Eq. 1b), while the latter is encapsulated by the lifetime in the trap (see $\tau_{\ominus}$ Eq. 1). We note that removing the lifetime restriction results in no clear localization of the highest atom numbers, as shown in Fig. B1 with an extended parameter space (compared to Fig. 2). The localization must otherwise be present as the trap lifetime becomes shorter for larger axial beam detuning imbalances and higher gradients in this considered space. We are therefore provided evidence that the dark-state loss is of fundamental importance for our observations.
We next compare the atom numbers. In the experiment, the highest atom number obtained is ${\sim}2.3\times 10^{5}$ atoms with the loading rate of $1.5\times 10^{6}$ atoms/s, which is expectedly bounded by that of the 2D MOT ($1.5\times 10^{9}$ atoms/s). A same-order-of-magnitude value for the 3D MOT loading has been reported in literature for a setup employing a push beam for comparable experimental parameters [45]; nonetheless, further optimizations of our loading rate can be made as discussed later (Secs. III.D and IV). By using the experimental flowrate in the analytical model’s Eq. 1 ($\mathcal{F}_{0}\times\mathcal{P}_{\oplus}\times\mathcal{P}_{\ominus}=1.5\times 10^{6}$ atoms/s), we obtain ${\sim}1.6\times 10^{5}$ atoms at the peak location (cross in Fig. 2), being in close agreement with the experiment (by $30\,\%$) and telling that our lifetime estimate reflects well the reality. Regarding the simulation, it predicts the maximum of ${\sim}3\times 10^{4}$ atoms, which  is  roughly  8  times
@@
less than in the experiment. This discrepancy is in small part due to the 2D MOT loading rate discrepancy ($13\,\%$; see Tab. 1) but mostly due to the loading being diminished by the choice of the radial beams offset (1 waist below the quadrupole field zero), as we report a simulated ${\sim}2.5$ times enhancement when a smaller offset is used (around a third of a beam waist), giving thus an overall ${\sim}3$ times difference between the simulation and experiment. This numerical atom number dependence on the offset is, however, different from the analytical result discussed below.
@@
Figure 3: The maximal 3D MOT atom number dependence on the radial beam offset, obtained using our analytical model (Sec. III.A). The vertical dashed line indicates the offset used in the experiment (by 1 radial beam waist below the quadrupole field zero). The normalization is done with respect to the result at this offset. The inset shows the top beam detuning (dashed-dotted line) and gradient (dotted line) dependence on this offset.
In Fig. 3, we provide an analytical calculation of the maximal 3D MOT atom number dependence on the radial beam offset. We first observe that by using the experimental beam offset (indicated by vertical dashed line) the atom number is seen to more than double compared to when no offset is used. Moreover, we observe that the atom number can be increased further by up to ${\sim}70\,\%$, given the beams are lowered by additional 2 radial beam waists and a larger axial beam detuning imbalance is used (see the inset) for a more downshifted $d_{\textsf{e}}$. Note that an approximate plateau is observed beyond this offset as the limit for the top beam detuning is reached (refer  to  Fig.
@@
B1), at which point the gradient has to decrease such that $d_{\textsf{e}}$ can continue to get lower. We expect the MOT to be empty for relatively large beam offsets, as the quadrupole MOT field needed for the confinement does not extend indefinitely, which is not considered in our analytical model. This detail is captured by our numerical model, which together with the included polarization effects can reasonably explain why the atom number should peak and eventually fall with increasing offset.
@@ C. The influence of metastable-state shel-
As implied by our analytical model (Sec. III.A), the 3D MOT loading efficiency critically depends on the dispersion the atoms accumulate (due to diffusion) during their transfer from the 2D MOT. This dispersion is statistically diminished by the metastable-state shelving (refer to the definition of $\sigma_{r}$ in Sec. III.A) but cannot be completely eliminated. It is thus natural to seek a configuration where the transfer speed is increased and the dispersion is kept low. We experimentally demonstrate this by misaligning the bottom beam, leading to a radiation-pressure imbalance without enhanced scattering. Using such a configuration, we report observing an increase in the experimental atom number (at the encircled point in Fig. 2) by roughly twice (to ${\sim}4.6\times 10^{5}$ atoms) with an approximately three-fold loading rate improvement (to ${\sim}4.5\times 10^{6}$ atoms/s). For this, the bottom beam misalignment of ${\sim}50$ mrad at the MOT center is used, with higher offset values resulting in an empty MOT. Moreover, we report that the radiation-pressure imbalance achieved using this special configuration makes it possible to trap with equally detuned axial beams, as our simulations also
confirm.
@@
Figure 4: Experimental data showing the 3D MOT atom numbers (normalized) for the cases where the bottom beam is misaligned (crosses) and aligned (circles) versus the repumping intensity $I_{tr\!f}$ in the transfer region (scaled by the saturation intensity $I_{sat,c}$). The misalignment is ${\sim}50$ mrad at the MOT center, and the remaining parameters are as in Fig. 2 for the maximal atom number. The normalization is done with respect to the aligned case for zero transfer-region repumping intensity. Note that the data is obtained in all cases with a constant intensity of repumping light present in a single retroreflected radial beam pair (see text).
@@
The misaligned configuration also opens a possibility of introducing repumping light in the transfer region, which is otherwise detrimental in the aligned configuration, as Fig. 4 implies. Here, we show experimentally obtained atom numbers in the 3D MOT for these respective cases (crosses and circles) versus the repumping intensity $I_{tr\!f}$ in the transfer region (overlapped with the top beam). We note that, for these measurements, repumping light of constant intensity is present in a retroreflected radial beam pair as in our original setting (see Sec. II.D). The respective increase and decrease versus $I_{tr\!f}$ serves as an experimental proof that the metastable-state shelving (to $5s5p\,^{3}\!P_{2}$) is essential for the 3D MOT loading dynamics. Indeed, with the misalignment the speed of the atoms is increased and results in a rapid transfer with less time to disperse outside the MOT capture, whereas for no misalignment the atoms are relatively slow and disperse in significant numbers before reaching the MOT region due to continuous heating. Moreover, with larger intensities, the atomic medium becomes increasingly saturated,  and
@@
hence the trends in Fig. 4 for the misaligned and aligned cases follow. Note that this figure implies that we can achieve an atom number improvement of ${\sim}2.5$ compared to the aligned configuration, for the former case using a relatively low repumping intensity in the transfer region ($I_{tr\!f}/I_{sat,c}\approx 0.2$). We can possibly enhance this number further by adding repumping light to the second pair of retroreflected radial beams, as the capture area for shelved atoms would be would be increased. In the next section, additional optimizations are discussed.
IV. Numerical optimization
We here demonstrate our simulation tool’s ability to improve the performance of an effusive oven experiment. We discuss how the 2D MOT loading can be enhanced by introducing changes to the 2D OM and the Zeeman slower. Also, we numerically verify the optimal angles for the radial beams in the 3D MOT (Sec. II.D).
Two important ways to manipulate the 2D OM performance include optimizing the beam power and the beam size (aperture and waist) [39]. The former is, however, resource intensive, considering the gains diminish rapidly due to the atom saturation while other parts of the setup (e.g., the 2D and 3D MOTs) may benefit more from the available laser power. In Fig. 5, we study the effects of increased beam power for different waist sizes at fixed aperture sizes. The data points connected with a straight line are obtained for the original aperture and beam waist (Tab. 1 case 4), while the data points connected with other kinds of lines are for a $4$ times larger aperture (requiring two-inch optics) and different beam waists (dashed for $1$ times the original, dashed-dotted for $2$, and dotted for $3$). We observe first that enlarging the aperture can result in an increased 2D MOT loading rate, depending on both the beam waist and the 2D OM power. An increase in this rate is a consequence of the atoms being exposed longer to the transverse cooling, provided the intensity is sizeable throughout the cooling region with the chosen beam waist. Indeed, as observed for lower intensities (below 30 mW), extending the waist can either be advantageous (dashed and dashed-dotted lines) due to effectively improved cooling, or detrimental (dotted line) due to effectively diminished cooling. For larger intensities, on the other hand, the saturation effect is present, such that wider waists result in the extension of a region of efficient cooling. In this case, it is particularly import-
ant to choose a waist that is not too narrow, so this effect can be taken full advantage of (observe the dotted line crossing the dashed one above 30 mW). Overall, the observations here agree with intuitive expectations, further validating the numerical model.
Figure 5: Numerical data showing the 2D MOT loading rate versus the 2D OM beam power (individual). Different aperture radii $a$ and beam waist radii $w$ of the 2D OM are considered. The aperature size ”$a\times 1$” and the waist size ”$w\times 1$” are from the original simulations, where $6$ mW of 2D OM beam power is also used (Tab. 1 case 4). The loading normalization is done with respect to this case.
Our ZS modeling (Eq. A8) has been derived with a purpose of tailoring a magnetic field to a given experimental environment, and we outline here how this can be achieved. With the initial oven conditions, ZS beam and 2D MOT parameters specified, one employs an algorithm that iteratively adjusts the slopes of the linear segments composing the field profile, with the aim of increasing the probability density of the atom speeds within the capture-speed range of the 2D MOT. To remove spurious fluctuations, each iteration step uses the same initial atom position and speed distributions with the diffusion (stemming from the radiation pressure; App. A1.B) turned off. In Fig. 6, we display the initial speed profile (light gray), and the final speed profile at the 2D MOT location (dark gray), which can be achieved in the iteration process. A concentrated probability density is seen to occur close to 0 m/s for the final profile (see the inset), corresponding to the atoms that have been captured by the 2D MOT. The tested field profile stems from our experiment where it more than quadruples  the  2D  MOT loading  rate (see
Sec. II.C).
Figure 6: Numerical data showing atom speed histograms before the Zeeman slowing (light gray) and at the 2D MOT location after the slowing (dark gray). The solid line displays the Maxwell-Boltzmann distribution for $440\,^{\circ}$ (as used in the original simulations; see Sec. II). The inset zooms in on the capturable speeds of the 2D MOT, showing a concentrated probability density near 0 m/s due to the trapping.
Finally, we verify for the 3D MOT side beams their optimal angles with respect to the radial axis perpendicular to the objectives axis. In Fig. 7, we display the numerical results for the 3D MOT atom numbers versus the beams angle (mirrored). As can be seen, the $25\,^{\circ}$ angle yields the highest number given our experimental constrains (dashed line limit). While larger angles can result in even higher numbers (see right of the dashed line), the beams may be clipped by our objectives [refer to the inset in Fig. 1(a)]. The optimum angle is found to be greater, at $30\,^{\circ}$, which can be explained to be a combination of a large repumping light volume in the science region (thus bringing many shelved atoms back into the cooling cycle) and the radial trapping forces being closer to perpendicular (thus resulting in more balanced forces). The resulting increase in the atom number is, however, marginal (by ${\sim}25\,\%$), and it is relatively robust against changes within a wide range of angles (from $20\,^{\circ}$ to $45\,^{\circ}$). Note that the findings here are not universal due to their dependence on the cell geometry. Implicitly, however, we are made aware that the optimum loading is in general not achieved at the standard $90\,^{\circ}$ angle between beam a-
xes (corresponding to $45\,^{\circ}$ in Fig. 7).
Figure 7: Numerical data showing the 3D MOT atom number versus the angle of the side beams mirrored with respect to the radial axis perpendicular to the objectives axis [refer to the inset in Fig. 1(a)]. The vertical dashed line indicates the limit beyond which our objectives may clip the beams. The normalization is done with respect to the point at the $25\,^{\circ}$ angle, as used in the experiment (see Sec. II). The remaining simulation parameters are the same as in Fig. 2, where the greatest atom number is obtained.
V. Conclusions
@@
This paper describes and analyzes an effusive oven setup integrating a moving molasses technique, where a 3D MOT is loaded using its axial beams creating a cold flux of atoms originating from a 2D MOT region. The loading is achieved through a detuning imbalance between the axial beams or a misalignment of the beam opposing the flux, with the combination of the two methods yielding atom number improvements (Sec. III.C). This technique completely excludes an additional push beam, commonly employed in effusive oven experiments, resulting in the 3D MOT being loaded in an advantageous laser beam configuration. Moreover, this provides a broad optical access to the scientific region of study, allowing for the realization of high NA optical tweezers and opening more laser-pathways for atom array manipulation, as highly desired in Rydberg-atom based simulation and computing [18, 19, 22, 23]. We find the metastable-state shelving and the dark-state loss to be essential for the MOT loading dynamics. Particularly for the dark-state loss, it is shown
by our analytical model to explain the existence of a well-localized parameter isle for highest atom numbers. This prediction is confirmed by our numerical model, which is more advanced than the analytical one (4-level versus 2-level model). The corresponding simulation tool is used to verify and optimize the loading to high accuracy. Notably, this tool simulates the full 3D atom dynamics starting at the hot source and ending at the 3D MOT, with the only free parameter being the atomic flowrate measured after the oven exit. Although the model is reliable for our purposes, including more detailed physics can enhance it. For instance, the initial velocity distribution, which affects the 2D OM performance and the downstream dynamics, may benefit from geometrical treatments tracking emerging particle trajectories [53, 57, 59] versus our treatment that attempts to average these trajectories. Moreover, in high optical depth regions encountered in, e.g., a ZS, collisions [60, 61] and multiple-scattering effects (attenuation and rescattering) can become of importance [62, 63]. The latter effects could be readily included from the model that we have been building upon [54], whereas the collisions could to first approximation be of a knock-out type within a van der Waals distance. Background collisions due to imperfect vacuum, on the other hand, could be modeled by introducing a vacuum-dependent decay time based on the main atom species considered [51, 64]. Additionally, collisions with the chamber walls may be required to be taken into account, depending on the surface binding energy of the species [56, 65]. With these improvements in mind, we envision that our model could be pivoted towards the development of a software for optimizations of arbitrary MOT loading experiments.
VI. Acknowledgments

### v2

@@ B. The experiment and comparison to
In Fig. 4, we compare the 3D MOT loading diagrams obtained analytically, experimentally, and numerically. The parameter space spans over different field gradients and top beam detunings. The shapes enveloping significant atom numbers are observed to take up roughly similar areas in all three cases, with the highest atom numbers (absolute values discussed below) obtained at the same data point locations in the experiment and simulation (dotted circle center, $48$ G/cm gradient at 10-10 MHz top beam detuning), while the analytical model prediction for the highest number is within the point’s neighborhood (cross, $44.8$ G/cm gradient at 7.3-7.3 MHz top beam detuning). From our analytical model, we find that its optimal atom number prediction occurs for the absolute transfer speed (from the 2D MOT) $|v_{max}|\approx$ 9 m/s, which leads to the transfer time $t\approx$ 36 ms. This time is approximately five times longer than the time $\tau$ until the atoms go dark as they shelve to the metastable state $5s5p\,^{3}\!P_{2}$ during their transfer to the 3D MOT, where they subsequently get revived due to the repumping in this region (Sec. II.D). Furthermore, we find the corresponding axial equilibrium position $d_{\textsf{e}}=-3$ mm (the minus sign refers to the position being below the 3D MOT quadrupole field zero), which is lower than the offset of the radial beams (2.4-2.4 mm, with the absolute being the waist size; see Sec. II.D). This difference is related to the atoms having an extended lifetime in the 3D MOT when not at the intensity maximum of the radial beams.
For the remaining shape differences in the diagrams of the experiment and simulation (more complex than analytical theory), this may be attributed to the simulation being sensitive to the 3D MOT magnetic field modeling, as we use an approximation to the quadrupole field created by the rectangular coils in our experiment (see Sec. II.D). The most significant magnetic field difference would be far from the 3D MOT center and in the transfer region itself, altering the loading dynamics.
We use the analytical model to explain the different features in the diagrams as follows:
@@
(i) For the dark area near the line where the axial beams are equally detuned (46-46 MHz; vertical dashed line), three parts are involved. Just to the left of this line, the atom transfer occurs in the opposite direction (larger top beam detuning). At this line, the transfer is obviously halted. Just to the right, the transfer is slow enough so the atoms disperse in significant numbers before reaching the 3D MOT region, with the higher gradients resulting in the dark area extending further towards resonance as the radial capture distance $d_{\textsf{c}}$ becomes smaller.
@@
(ii) The dark area on the right side of the diagrams (delineated by the slanted dashed line) is, on the other hand, obtained due to the axial equilibrium position $d_{\textsf{e}}$ being far enough below the radial 3D MOT beams center so the confinement is significantly diminished, or even non existent due to the radial beams not encompassing $d_{\textsf{e}}$. The boundary appears linear as $d_{\textsf{e}}$ is proportional to the axial beam detuning difference while inversely proportional to the gradient.
@@
(iii) Regarding the bright isle with the highest atom numbers, its location depends on the MOT confinement and the dark-state loss. The former expresses itself in the capture from the atomic  beam (see the error functions in Eq. 1) and the relative positioning of the radial 3D MOT beams (see the Gaussian in Eq. 2), while the latter is encapsulated by the decay of the trap (see $\Gamma_{\text{loss}}$ in Eq. 3). We note that removing the decay results in the absence of a clear localization of the highest atom numbers, as shown in Fig. 7 with an extended parameter space (compared to Fig. 4). The localization must otherwise be present as the trap lifetime becomes shorter for larger axial beam detuning imbalances and higher gradients in this considered space. We are therefore provided evidence that the dark-state loss is of fundamental importance for our observations.
We next compare the atom numbers. In the experiment, the highest atom number obtained is ${\sim}2.3\times 10^{5}$ atoms with the loading rate of $1.5\times 10^{6}$ atoms/s, which is expectedly bounded by that of the 2D MOT ($1.5\times 10^{9}$ atoms/s). A same-order-of-magnitude value for the 3D MOT loading has been reported in the literature for a se-
@@
tup employing a push beam for comparable experimental parameters [45]. Nonetheless, there are adjustments to be made to greatly increase our loading rate, and, in Sec. III.C, we discuss the numbers that can be attained. By using the experimental flow rate in the analytical model’s Eq. 3 ($\mathcal{P}\times\mathcal{F}_{0}=1.5\times 10^{6}$ atoms/s), we obtain ${\sim}1.6\times 10^{5}$ atoms at the peak location (cross in Fig. 4), which is in close agreement with the experiment (by $30\,\%$) and telling that our lifetime estimate reflects the reality well. Regarding the simulation, it predicts the maximum of ${\sim}3\times 10^{4}$ atoms, which is roughly 8 times less than in the experiment. This discrepancy is in small part due to the 2D MOT loading rate discrepancy ($13\,\%$; see Tab. 1) but mostly due to the loading being diminished by the choice of the radial beams offset (1 waist below the quadrupole field zero), as we report a simulated ${\sim}2.5$ times enhancement when a smaller offset is used (around a third of a beam waist), giving thus an overall ${\sim}3$ times difference between the simulation and experiment. This numerical atom number dependence on the offset is, however, different from the analytical result discussed below.
@@
Figure 5: The maximal 3D MOT atom number dependence on the radial beam offset, obtained using our analytical model (Sec. III.A). The vertical dashed line indicates the offset used in the experiment (by 1 radial beam waist below the quadrupole field zero). The normalization is done with respect to the result at this offset. The inset shows the top beam detuning (dash-dotted line) and gradient (dotted line) dependence on this offset.
In Fig. 5, we provide an analytical calculation of the maximal 3D MOT atom number dependence on the radial beam offset. We first observe that by using the experimental beam offset (indicated by the vertical dashed line)
@@
the atom number is seen to more than double compared to when no offset is used. Moreover, we observe that the atom number can be increased further by up to ${\sim}70\,\%$, given the beams are lowered by an additional 2 radial beam waists and a larger axial beam detuning imbalance is used (see the inset) for a more downshifted $d_{\textsf{e}}$. Note that an approximate plateau is observed beyond this offset as the limit for the top beam detuning is reached (refer to Fig. 7), at which point the gradient has to decrease such that $d_{\textsf{e}}$ can continue to lower. We expect the MOT to be empty for relatively large beam offsets, as the quadrupole MOT field needed for the confinement does not extend indefinitely, which is not considered in our analytical model. This detail is captured by our numerical model, which together with the included polarization effects can reasonably explain why the atom number should peak and eventually fall with increasing offset.
@@ C. The influence of metastable-state shel-
As implied by our analytical model (Sec. III.A), the 3D MOT loading efficiency critically depends on the dispersion the atoms accumulate (due to diffusion) during their transfer from the 2D MOT. This dispersion is statistically diminished by the metastable-state shelving (refer to the definition of $\sigma_{r}$ in Sec. III.A) but cannot be completely eliminated. It is thus natural to seek a configuration where the transfer speed is increased and the dispersion is kept low. We experimentally demonstrate this by misaligning the bottom beam (see Fig. 1), leading to a radiation-pressure imbalance without enhanced scattering. Using such a configuration, we report observing an approximately 2 times increase in the experimental atom number (to ${\sim}4.6\times 10^{5}$ atoms). For this, the bottom beam misalignment of ${\sim}50$ mrad at the MOT center is used, with higher offset values resulting in an empty MOT. Moreover, we report that the radiation-pressure imbalance achieved using this special configuration makes it possible to trap with equally detuned axial beams, as our simulations also confirm. Note that, in the further discussion, we focus only on the misaligned case having the same MOT parameters (including the axial beam de-
tuning imbalance) as in the aligned case.
@@
Figure 6: Experimental data showing the 3D MOT atom numbers (normalized) for the cases where the bottom beam is misaligned (crosses) and aligned (circles) vs the repumping intensity $I_{tr\!f}$ in the transfer region through the top beam (scaling with the saturation intensity $I_{sat,c}$). The misalignment is ${\sim}50$ mrad at the MOT center, and the remaining parameters are as in Fig. 4 for the maximal atom number. The normalization is done with respect to the aligned case for zero transfer-region repumping intensity. Note that the data are obtained in all cases with a constant intensity of repumping light present in a single retroreflected radial beam pair (see text).
@@
The misaligned configuration also opens a possibility of introducing repumping light in the transfer region, which is otherwise detrimental in the aligned configuration, as Fig. 6 implies. Here, we show experimentally obtained atom numbers in the 3D MOT for these respective cases (crosses and circles) versus the repumping intensity $I_{tr\!f}$ in the transfer region (overlapped with the top beam). We note that, for these measurements, repumping light of constant intensity is present in a retroreflected radial beam pair as in our original setting (see Sec. II.D). The respective increase and decrease versus $I_{tr\!f}$ serves as an experimental proof that the metastable-state shelving (to $5s5p\,^{3}\!P_{2}$) is essential for the 3D MOT loading dynamics. Indeed, with the misalignment the speed of the atoms is increased and results in a rapid transfer with less time to disperse outside the MOT capture, whereas for no misalignment the atoms are relatively slow and disperse in significant numbers before reaching the MOT region due to heating. Adding repumping to the transfer region extends the time the atoms are affected by the blue light, creating (i) for the misaligned case, further acceleration downwards in the region where the top and bottom beams do not overlap, and (ii) for the aligned case, further detrimental heating leading to radial  dispersion. With  larger
@@
intensities, the atomic medium becomes increasingly saturated, and hence the trends in Fig. 6 for the misaligned and aligned cases follow. Note that this figure implies that we can achieve an atom number improvement of ${\sim}2.5$ compared to the aligned configuration, for the former case using a relatively low repumping intensity in the transfer region ($I_{tr\!f}/I_{sat,c}\approx 0.2$). We can possibly enhance this number further by adding repumping light to the second pair of retroreflected radial beams, as the capture area for shelved atoms would be increased.
There exist adjustments in our system that would add significantly to the loading enhancement. One of the adjustments concerns the oven temperature (Sec. II.A), while the other concerns the ZS performance (Tab. 1). Our oven temperature of $440\,^{\circ}$C is relatively modest compared to the temperature of, e.g., $550\,^{\circ}$C reported in the literature [39, 42]; assuming that the capturable flux increases as $P_{\text{Sr}}(T)\times T^{-2}$ $P_{\text{Sr}}(T)$ is a semi empirical equation for vapor-pressure of strontium at temperature $T$ in Kelvin [62, 63], and $T^{-2}$ is the scaling of the 1D Maxwell-Boltzmann distribution in the low-speed limit], we expect around 25 times loading enhancement. By optimizing the ZS stage, we expect to bring the enhancement to two orders of magnitude [39, 42]; installing servomotors for the optimization may be a good practical choice [64], although we note that it is costly compared to using a numerical algorithm [42, 49]. All in all, by increasing the oven temperature to $550\,^{\circ}$C, optimizing the ZS stage, and working in the misaligned configuration while employing repumping considerations above, we expect the 3D MOT loading rate to exceed $4\times 10^{8}$ atoms/s and be within an order of magnitude with the best performing Sr loading experiments [38, 39, 42, 65].
IV. Conclusions
@@
This paper describes and analyzes a setup integrating a moving molasses technique, where a 3D MOT is loaded using its axial beams creating a cold beam of atoms from the location of a 2D MOT (Fig. 1). The highest loading is achieved by misaligning the bottom axial beam with a different detuning than the top beam (by ${\sim}1\,\Gamma$). Loading rates exceeding $4\times 10^{8}$ atoms/s are achievable with a high oven temperature ($550\,^{\circ}$C) and an optimized ZS, being well sufficient for obtaining dense red MOTs and large Sr tweezer arrays [16, 23, 52, 66]. This technique completely excludes an additional push beam and in-vacuum parts, while resulting in the 3D MOT being loaded in an advantageous laser beam configuration. Moreover, it provides a broad optical access to the scientific region of study, allowing for the realization of high  NA  optical  tweezers
and opening more laser pathways for atom array manipulation, as highly desired in Rydberg-atom based simulation and computing [18, 19, 22, 23].
We find the metastable-state shelving and the dark-state loss to be essential for the MOT loading dynamics. Particularly for the dark-state loss, it is shown by our analytical model to explain the existence of a well-localized parameter isle for highest atom numbers. This prediction is confirmed by our numerical model, which is more advanced than the analytical one (four-level versus two-level model). The corresponding simulation tool is used to verify the loading to high accuracy and can be employed in the design and optimization of effusive oven experiments. Indeed, this tool simulates the full 3D atom dynamics starting at the hot source and ending at the 3D MOT (with the only free parameter being the atomic flow rate measured after the oven exit) and allows for arbitrary beam geometries while taking into account the effects of the magnetic fields and the Poissonian atom decay.
As the final remark, we report employing our moving molasses technique in a separate experiment with alkali atoms (cesium) [67] to load large tweezer arrays, thus demonstrating that this technique can be extended beyond alkaline earth metals.
V. Acknowledgments